OpenStudy (anonymous):

Find the greatest common factor for the group of terms. -66a^3, 11a^5

myininaya (myininaya):

-66a^3=-6*11*a*a*a 11a^5=11*a*a*a*a*a they have an 11 in common they also have a*a*a=a^3 in common so the gcf is 11a^3

OpenStudy (anonymous):

Thanks